OpenStudy (anonymous):

i have a question that i need help solving. Its moments and centers of mass. The question states that the region bounded by the parabola y=x^2 and the line y=4

OpenStudy (anonymous):

find the center of mass of a thin plate of constant density and covering given the region and its thin plates with constant density

OpenStudy (anonymous):

have you started working on this?

OpenStudy (anonymous):

yes trying to i kno that bar x =0 and bar y =12/5, but i cant find the bar y

OpenStudy (anonymous):

nvm got it

OpenStudy (anonymous):

nice :)

OpenStudy (anonymous):

nvm need help

OpenStudy (anonymous):

and in order to find bar y, i have to to find Mx, which to find bar y, you have to do Mx/m which m =-4/3

OpenStudy (mathmate):

@mrGrimm

OpenStudy (mathmate):

@mrGrimm  This is how you can set up the double integral. The area A is obtained by \(\int\int dx~dy \) over the region. Mx is obtained by \(\int\int ~y~ dx~dy \) over the region. The limits have been shown in the above figure, namely first integrate along x, which means that the limits must be expressed in term of y, giving [-sqrt(y), +sqrt(y)]. The second integrate is along y, so [0,4]. Finally, \(\bar{y}\) = \(\dfrac{M_x}{A}\), as you probably know already. 12/5 is the correct answer, as you stated.